OpenStudy (anonymous):

Can someone explain to me simultaneous equations x + y = 8 and y - x = -2, then y =

OpenStudy (amistre64):

the solution has to fit both conditions

OpenStudy (amistre64):

x + y = 8 let x = (4+n), let y = (4-n) (4+n) + (4-n) = 8 fits this one solve in y-x = -2 (4-n) - (4+n) = -2

OpenStudy (anonymous):

wait what? this is for substitution method or elimination method?

OpenStudy (espex):

@amistre64's method is neither method. Given that you have no leading coefficient on either variable, you can simply use substitution method.

OpenStudy (amistre64):

its a substitution method, something Diophantus might do.